### v1

# Literature Filtering for Systematic Reviews with Transformers

Literature Filtering for Systematic Reviews with Transformers
John Hawkins $0$-0001-6507-3671 Transitional AI Research GroupSydenyAustralia43017-6221  and  David Tivey Royal Australasian College of Surgeons (RACS)Australia
Systematic Review, Literature Screening, Machine Learning, BERT, Transfer Learning, Natural Language Processing
1. Abstract
Identifying critical research within the growing body of academic work is an essential element of quality research. Systematic review processes, used in evidence-based medicine, formalise this as a procedure that must be followed in a research program. However, it comes with an increasing burden in terms of the time required to identify the important articles of research for a given topic. In this work, we develop a method for building a general-purpose filtering system that matches a research question, posed as a natural language description of the required content, against a candidate set of articles obtained via the application of broad search terms. Our results demonstrate that transformer models, pre-trained on biomedical literature then fine tuned for the specific task, offer a promising solution to this problem. The model can remove large volumes of irrelevant articles for most research questions.
2. Introduction
Systematic Reviews are an essential aspect of evidence-based medicine. They ensure that researchers are aware of the most relevant literature on a particular topic. They are also time-consuming, error-prone and costly to undertake(Dinter2021b, ). The expanding volumes of literature and growing number of active research areas are increasing the burden of thorough systematic reviews(Bastian2010, ; Miwa2014, ).
The problem of matching a research question to a set of articles sits within the broader problem of matching queries to documents, which is plagued by the issue of algorithmically identifying what is semantically relevant between a query and document pair(Ostendorff2020, ). A common approach to building software applications that assist the systematic review process focuses on tools that refine database literature searches and results, including the use of empirical methods for refining search terms (Wilczynski2004, ; Zdravevski2019, ; Weier2020, ), automated query generation(Scells2020, ), and recommendation methods based on the researcher’s publication profile(Chen2016, ) and machine learning models that emulated the identification of articles of clinical importance(Aphinyanaphongs2005, ).
Recent development has focused on methods to improve the efficiency of the initial literature screening, including processes that reduce the human workload through question-specific predictive models built on an initial set of human reviewer labels(Frunza2010, ; Bannach-Brown2019, ), expanded to include other kinds of meta-data(Olorisade2019, ) and applications that update the prioritisation of abstracts in an iterative feedback process (Jonnalagadda2013, ; Miwa2014, ; Kontonatsios2020, ; Bravo2021, ), or expansion of the articles for consideration using a similarity space and initial human selected abstracts (Pham2021, ). Recent machine learning studies for automation of systematic reviews focus on alternative machine learning approaches to building these task/question specific models, with Convolutional Neural Networks (CNNs) (Dinter2021a, ), and Transformer Networks (Aum2021, ). Multiple studies are making incremental steps toward more general models that reuse what has been learned from previous systematic reviews in different kinds of transfer learning. This includes learning document embeddings that capture information relevant to literature filtering(Kontonatsios2020, ), and clustering candidate article abstracts with the articles from previous literature reviews to identify high-quality candidates(Cawley2020, ).
However, there is some question as to whether these tools provide any real value to systematic review teams(Gartlehner2019, ). The tools Abstrackr(Wallace2012, ) and EPPI-Reviewer(Thomas2010, ) can both be trained on custom datasets to filter articles toward a subset for human reviewers. Independent evaluation has indicated that these tools will reduce the burden on human reviewers by a maximum of 50-60%, respectively. However, on some research questions, they perform considerably worse(Tsou2020, ). Arguably, the biggest challenge for this approach is the requirement for an initial set of labelled data for the specific question that is large enough to provide a viable model(Aum2021, ).
Alternative approaches use specialised, but general-purpose models, to identify factors like the presence of randomised control trials within a given study(Cohen2015, ), or the specific text sections that correspond to the elements of the PICO(Richardson1995, ) framework (population; intervention; comparator; outcomes) (Wallace2016, ; Kang2019, ; Jin2020, ). Multiple factors about a paper can be used to filter articles in a systematic review, including the scientific rigor of the content (Ambalavanan2020, ), the quality of the research irrespective of the content(Lokker2023, ). There are multiple datasets available for these tasks(Nye2018, ; DeYoung2020, ; Ambalavanan2020, ), and multiple modelling approaches applied to this task with a recent focus on the use of neural networks from Long Short Term Memory (LSTM) (Jin2018, ) to Transformer models(Schmidt2020, ; Ambalavanan2020, ).
Review articles from the systematic review literature itself have concluded that machine learning based filters that reduce the volume of articles needed to be reviewed are the most promising line of research in the automation of systematic reviews (Marshall2019, ; OMara-Eve2015, ).
The Transformer model(Vaswani2017, ) has proven to be an effective tool for building text-based classification systems across a range of domains. Part of the utility comes from the fact that it can be pre-trained on the language from a domain and then used for specific tasks, allowing the previous learning to be ’transferred’ and applied. It has been shown that transformer models pre-trained on academic abstracts can be used to create state-of-the-art classifiers for systematic review systems, but only when they train the models on data for a specific task (Aum2021, ), extending previous work with feed-forward neural networks to learn a document embedding that can be employed via transfer learning for a specific systematic review task(Kontonatsios2020, ). The requirement for task-specific training places a burden on teams in terms of the time and human effort required to provide labels for the initial classifier. The open research question is whether we can use modern natural language processing transformers to develop general-purpose filters that learn the general correspondence between the phrasing of a research question and the relevant articles.
This research outlines a dataset and comparison of methods for a general-purpose systematic review model. A single model is trained that is given input data combining a research question and information about a specific article. It then produces a score indicating the likelihood that the specific article would be included by a human reviewer during a literature pre-screen. Our results demonstrate that using BioBERT(Lee2020, ) version of the BERT Transformer(Devlin2019, ) pre-trained on biomedical literature can contribute a promising solution to this problem, but the ideal approach depends on the intended process for using a model to improve the systematic review process. We discuss the observed relationships between research questions and machine learning approaches, and conclude that more work is needed to determine the ideal model for all systematic reviews.
3. Methodology
Figure 1. Schematic representation of the standard machine learning approach to building a systematic review filter tool. 1. Research question is proposed, 2. Database search is conducted, 3. List of candidate articles is compiled, 4. Human annotators review a small selection of articles, 5. Small annotated collection is used to train a machine learning filter, 6. The machine learning is applied to the remaining documents, 7. Producing the final set for complete review by human experts.
The standard approach to building systematic review tools with machine learning shown in Figure 1. The standard model requires an initial labelled set to seed the filter model.
Figure 2. Schematic representation of the general systematic review filter process. The first five steps define how the general filter is trained, the last two the application to new research questions. 1. A set of training research question are taken, 2. Database searches are conducted for each, 3. Lists of candidate articles are compiled, 4. Human annotators review these training questions to create an annotated set, 5. The general filter is trained. 6. The machine learning is applied to filter documents for a new query, 7. Producing the final set for complete review by human experts.
Our alternative approach is depicted schematically in Figure 2. The model is trained on a new dataset derived from multiple research questions with human-annotated lists of articles returned from a database search. The bottom frame demonstrates that this general purpose filter can now be applied to future research questions without the need for human annotators to provide the initial labeled set. Our approach learns an abstract relationship between research questions and articles that transcends individual research questions.
3.1. Dataset Curation
Table 1. Medical Research Questions Datasets
Project Clinical Area Description Records Relevant Incl Rate
LANB Drug: analgesia Post-surgical Nerve Blockade 766 314 41.0%
PID Drug: Antibody infusion Primary Immunodeficiency 12269 1052 8.6%
CIDP Drug: Antibody infusion Neurological Inflammation 4888 634 13.0%
PBRT Radiotherapy Partial Breast Radiotherapy 624 154 24.7%
IORT Radiotherapy Intraoperative Radiotherapy 1873 93 5.0%
ESG Surgery Endoscopic Sleeve Gastroplasty 4109 314 7.6%
EMR Surgery Endoscopic Mucosal Resection 3759 622 16.5%
CARGEL Surgery Ligament Substitution 3626 66 1.8%
LMTA Surgery Lung Microwave Tissue Ablation 452 68 15.0%
VERT Surgery Oesteoporotic Vertebral Fractures 2359 643 27.3%
We restrict our initial analysis to research questions that are concerned with Therapeutics, a constrained domain within the scope of medical research topics. We identified 10 research questions from the ASERNIP-S systematic review archive for which a team of human researchers had manually labelled thousands of articles as either relevant or not relevant. These initial questions are summarised in Table 1, showing the total number of abstracts retrieved from bibliographic database (e.g, PubMed) searches and the number deemed relevant by the human reviewers.
Each of these therapeutics questions has a corresponding research question file, which is a plain text description of the research program. Each research question was derived from the documentation that formed the source for both the construction of the database query to extract the candidate articles and the guidelines by which the human reviewers identified relevant research.
In addition to the natural language expression of the overall research question, we also generated natural language expressions for each of the four PICO elements independently. The PICO framework has been shown to be well suited to representing therapy questions(Huang2006, ), and we conduct experiments to determine whether text based machine learning tools are able to exploit this alternative framing of the research question.
The total dataset contains $34,725$ records with an overall inclusion rate of $11\%$. However, as shown in Table 1 the inclusion rate varies significantly between the research questions, as does the volume of articles returned from the database searches.
3.2. Machine Learning Models
Table 2. Text Match Feature Functions
Feature Function Description Library
Length Length of text block in chars -
WC Count of words in text block -
WL Length of words in text (max and mean) -
CWD Proportion of words that are not stop words -
Match Prop Proportion of question key words in article text -
JD Jaro distance between question and article text Jellyfish
LD Levenshtein distance between question and article text Jellyfish
JI Jaccard index between question and article text Textdistance
SD Sorensen distance between question and article text Textdistance
TFIDF Term Frequency Inverse Document Frequency Scikit-Learn
We approach this problem using four general machine learning strategies. The first strategy is a baseline approach using a set of standard feature engineering approaches and general-purpose machine learning algorithms ranging from simple to complex. In this approach the research question is contained in a single text field of the data and compared with the article’s title and abstract using text matching algorithms. Each of these text matching approaches generates a numeric value which either describes the context of the research article (title and abstract) or indicates the similarity of the text in the question and the article (title and abstract). The text features and matching approaches are drawn from multiple python packages and compiled into the texturizer python package (texturizer, ); the full set of feature generating functions we use are summarised in Table 2. Note, that the TFIDF approach is only used in some of the experiments, and in all cases this is indicated in the descriptions of the results. The majority of the algorithms used are drawn from the SciKit Learn package (full details of which can be found in the code base).
In the second approach, we extend the baseline models to include the research question text posed as the four independent PICO framework components. Each of the PICO components are matched to the article title and abstract independently. The effect of which is to quadruple the text match features, but to give the model an indication of what the specific match is between research question and article. For example, one article might have a high match on population and intervention, whereas another might have a high match of intervention and outcome.
In the third approach, we employ a set of pre-trained BERT Transformer models (Devlin2019, ) to generate sentence similarity scores as new features. We continue to use the standard text match features from strategies one and two, but simply append the BERT similarity features to the dataset. This approach is done using both the standard natural language representation of the research question and the PICO framework text. We use two general versions of BERT, the BioBERT model(Lee2020, ) and the BlueBERT model(peng2019, ). Summary details for these pre-trained BERT models is shown in Table 3, including the number of heads, layers and the vocabulary size of each model. These parameters provide an approximation to the complexity of each model.
In the final approach, we fine-tune the pre-trained BERT Transformer models to directly predict whether an article should be included for full-text review by researchers. In this final approach, no other features are being used, the BERT model is fed the research question natural language text, followed by the text block end token and then the article’s title and abstract concatenated together. This input representation involves reuse of one of the pre-training pathways for BERT models in which they are fed two sentences and trained to predict if the sentences have the same meaning. In our application, the BERT model is trained to predict if there is sufficient match between the research question and the article to warrant a full-text review.
Table 3. Pre-Trained Transformers
Transformer Model Heads Layers Vocab Size Tuning Data
BioBERT 12 12 28996 PubMed
BlueBERT 12 12 30522 PubMed + MIMIC
BlueBERTxL 16 24 30522 PubMed + MIMIC
3.3. Experimental Design
We aim to design a system that can identify relevant abstracts based on a new research question proposed for systematic review. The ideal system cannot rely on a bootstrap set of initial labelled examples for the specific research question. To align our testing protocol with this goal we use a variation of the cross-validation methodology in which each fold of the cross-validation consists of all data for one research question. This Leave-One-Question-Out (LOQO) method provides a close approximation to real-world usage, we evaluate how well the model can identify the relevant articles based purely on the correspondence between the content of the research question and the content of the article’s title and abstract.
3.4. Evaluation Metrics
To evaluate our results we use multiple different metrics. The Area Under the Curve ($AUC$)(Hanley1982, ) provides a value bounded between zero and one that captures the overall ability of a model to rank all articles(Melo2013, ). In addition, we use two metrics to evaluate the performance of the model at the upper and lower end of the article ranking, each of which cater to specific use cases of the models. We first consider the use case of eliminating irrelevant candidate articles (a dominant method by which machine learning filtering tools can provide efficiency for systematic review processes(Marshall2019, ; OMara-Eve2015, )). To evaluate models for this application we calculate the accuracy in the bottom 50% ($AccBot50\%$). We take the 50% lowest scoring records for a research question and then calculate the percentage of them that were annotated as irrelevant by the human reviewers. The calculation is shown in Equation 3.1:
[…]
Where $n$ is the number of articles returned for question $Q$. The function $f(Q_{i})$ returns the model score for article $i$ of question $Q$, and the function $l(Q_{i})$ returns the label.
4. Results
We show the results for both standard machine learning approaches and transformer models in Table 4 and Table 5. These table display the $AUC$, and $AccBot50\%$, respectively. Each modelling approach is displayed as a row with the results for each of the research questions in the corresponding column. We have highlighted in yellow the best (or equal best) score for each research question.
Table 4. AUC by Model and Research Question
experiment CARGEL CIDP EMR ESG IORT LANB LMTA PBRT VERT PID
1 - Linear 0.436 0.438 0.52 0.612 0.533 0.536 0.57 0.409 0.503 0.506
1 - Naive Bayes (GNB) 0.694 0.611 0.696 0.5 0.78 0.526 0.614 0.678 0.639 0.745
1 - Naive Bayes (CNB) 0.584 0.587 0.625 0.659 0.605 0.52 0.524 0.563 0.646 0.52
1 - SVM Classifier 0.665 0.493 0.646 0.613 0.594 0.549 0.631 0.583 0.591 0.683
1 - ExtaTrees 0.734 0.568 0.654 0.611 0.758 0.503 0.573 0.644 0.581 0.637
1 - TFIDF-ExtaTrees 0.829 0.769 0.77 0.674 0.748 0.535 0.614 0.696 0.591 0.837
1 - LGBM-TFID 0.834 0.819 0.795 0.685 0.812 0.537 0.658 0.697 0.555 0.841
2 - Linear-PICO 0.505 0.452 0.402 0.507 0.502 0.515 0.56 0.393 0.466 0.725
2 - ExtaTrees-PICO 0.659 0.621 0.742 0.606 0.712 0.499 0.596 0.671 0.616 0.763
2 - TFIDF-ExtaTrees-PICO 0.847 0.755 0.75 0.634 0.74 0.529 0.591 0.699 0.62 0.853
2 - LGBM-TFID-PICO 0.857 0.764 0.798 0.694 0.721 0.562 0.653 0.706 0.598 0.855
3 - LGBM-BioBERT 0.841 0.818 0.805 0.696 0.811 0.535 0.67 0.696 0.554 0.831
3 - LGBM-BlueBERT 0.845 0.816 0.792 0.689 0.824 0.539 0.647 0.687 0.556 0.838
3 - LGBM-BlueBERTxL 0.832 0.82 0.809 0.677 0.812 0.547 0.657 0.703 0.575 0.835
3 - LGBM-BlueBERT-PICO 0.865 0.76 0.794 0.702 0.735 0.568 0.63 0.706 0.578 0.852
4 - Fine-tune-01e-BioBERT 0.814 0.496 0.502 0.806 0.808 0.79 0.79 0.501 0.81 0.498
4 - Fine-tune-02e-BioBERT 0.873 0.5 0.503 0.497 0.626 0.532 0.777 0.502 0.505 0.685
The AUC describes the ability of the model to rank all articles for the given research question, a value of one would indicate perfect ranking. In some instances we see values very close to 0.5 in Table 4 indicating that in some instances the models are performing no better than chance. Note that all scores were generated using the LOQO process to show the performance of truly novel questions and a list of candidate articles.
Table 5. Bottom 50% Accuracy by Model and Research Question
experiment CARGEL CIDP EMR ESG IORT LANB LMTA PBRT VERT PID
1 - Linear 0.977 0.855 0.842 0.944 0.956 0.619 0.885 0.692 0.718 0.915
1 - Naive Bayes (GNB) 0.99 0.903 0.916 0.848 0.988 0.608 0.889 0.853 0.808 0.964
1 - Naive Bayes (CNB) 0.991 0.899 0.892 0.956 0.96 0.674 0.832 0.571 0.824 0.912
1 - SVM Classifier 0.99 0.872 0.897 0.945 0.962 0.619 0.907 0.808 0.778 0.951
1 - ExtaTrees 0.993 0.893 0.902 0.945 0.983 0.587 0.867 0.827 0.772 0.945
1 - TFIDF-ExtaTrees 0.996 0.952 0.945 0.955 0.991 0.616 0.898 0.862 0.78 0.981
1 - TFIDF-LGBM 0.997 0.97 0.954 0.959 0.994 0.611 0.912 0.849 0.761 0.981
2 - Linear-PICO 0.982 0.857 0.794 0.927 0.953 0.598 0.885 0.696 0.706 0.966
2 - ExtaTrees-PICO 0.99 0.907 0.931 0.944 0.982 0.595 0.894 0.837 0.796 0.968
2 - TFIDF-ExtaTrees-PICO 0.997 0.947 0.938 0.951 0.987 0.606 0.889 0.853 0.803 0.984
2 - TFIDF-LGBM-PICO 0.997 0.955 0.951 0.96 0.978 0.621 0.907 0.849 0.786 0.984
3 - LGBM-BioBERT 0.994 0.965 0.955 0.962 0.993 0.621 0.92 0.843 0.768 0.978
3 - LGBM-BlueBERT 0.997 0.965 0.952 0.96 0.993 0.619 0.907 0.827 0.759 0.982
3 - LGBM-BlueBERTxL 0.996 0.97 0.957 0.959 0.994 0.621 0.912 0.853 0.774 0.979
3 - LGBM-BlueBERT-PICO 0.997 0.95 0.949 0.961 0.98 0.627 0.907 0.862 0.783 0.984
4 - Fine-tune-01e-BioBERT 0.97 0.884 0.886 0.971 0.963 0.958 0.952 0.887 0.964 0.887
4 - Fine-tune-02e-BioBERT 0.982 0.887 0.886 0.885 0.912 0.891 0.948 0.887 0.888 0.955
Looking at the results in tables 4 and 5, the best-performing models tend to be toward the bottom of the table where we the modelling approaches make use of the pre-trained transformer models. There does not appear to be a single modelling approach that brings us toward optimal performance across all research questions, however, the fine-tuning approach resulted in improved performance across most research questions. Furthermore, this approach was the only approach to deliver functional AUC for the LANB and VERT questions. All other approaches provided performance that was barely better than random for these questions.
Finally, even within a given category of modelling approach (our 4 experiments), there is not a single clear winner in most instances. For example, even though fine-tuning a pre-trained transformer is the strongest overall modelling approach, the performance over one or two epochs of training varies considerably between research questions. This makes it difficult to determine which model will perform optimally for a future research question.
5. Discussion
We conducted a general exploration of machine learning approaches to filtering articles returned in a bibliographic database search to identify evidence for automated inclusion or exclusion. We have conducted these experiments using a rigorous Leave-One-Question-Out methodology, which allows us to understand the ability of each approach to perform when presented with truly novel research questions.
Our experiments provides substantial evidence that the use of pre-trained transformer models offers promise for improving systematic review efficiency through general-purpose machine learning filters. After fine tuning, these models can generalise to recognise articles that are inappropriate for inclusion in the systematic review of genuinely novel research questions. When considering the use case of filtering returned articles to exclude the bottom 50%, the highest accuracy was achieved in 5 of the 10 questions by fine tuning the BioBERT model for a single epoch of training. In the remaining 5 questions, we obtained the highest (or equal highest) accuracy using the pre-trained BERT models to generate similarity match scores between the research questions and article abstract.
The ten research questions are shown again in Table 6 with annotation to define key properties of these inquiries for both the PICO element and standard question formats. These questions were authored by multiple researchers, resulting in various language styles. As such, they reflect a real-world test for automated screening and the flexibility of Natural Language Processing algorithms to identify relevant articles in a dataset extracted from bibliographic databases. Before starting these experiments we hypothesised that questions with greater clinical area coverage would perform better, this was not observed; indicating that the therapy aspect of the research question is more important.
Using Microsoft Word, readability statistics were generated for the standard format questions. Except for the LANB and VERT questions, all questions had a Flesch-Kincaid Grade Level (FKGL)(Kincaid1975, ) above the average level of 15.87 for journal articles, as reported by Kandula and Zeng-Treitlar (2008)(Kandula2008, ). As such, the language used by researchers when drafting questions aligns with a journal article corpus. The low FKGL scores may account for the general poor performance of the NLP models, except for Fine-tuned BioBERT, when analysing the LANB and VERT questions.
In examining the question characteristics and the four questions (PID, CIDP, PBRT, and EMR) that returned the lowest performance in the Fine-tuned BioBERT model (AUC approx 0.5), the incorporation of the Comparator and Outcome PICO components had no substantial impact on the AUC results. Interestingly, two of the four questions with the lowest AUC (Fine-tuned BioBERT) possessed a comprehensive PICO representation (CIDP and EMR), while only the PI elements were present in the PID and PBRT questions.
In experiment 4, four questions returned sub-optimal AUC results for the Fine-tuned BioBERT model (PID, CIDP, PBRT and EMR). Two of these, as discussed above, performed well with a PICO representation. The remaining two, performed best using transfer learning with a larger BlueBERT model. This suggests that some research questions contain sufficient complexity (or subtlety) that demands a larger model with a large vocabulary in order to recognise a match between question and article.
Table 6. Comparison of Research Question Properties
PICO Format Standard Question
Question Population Intervention Comparator Outcome PICO Content Exclns Words Flesch-Kincaid GL
LANB B-DS N-DS SC DS-MO PI_O Low No 46 12.9
PID F-DS N-MS SC DS-MO PI__ Low No 21 25.1
CIDP B-DS N N-MS-EX DS-MO PICO Med Yes 112 19.2
PBRT ~ N-MS N-MS DS-MO PI__ Low Yes 29 20.5
IORT F-KW N-MS N DS-MO PIC_ Med No 72 21.5
ESG F-DS N-DS N-DS DS-MO PICO High No 304 24.2
EMR F-DS N-DS N-DS DS-MO PICO High No 150 23.0
CARGEL F-KW-MS N SC DS-MO PI__ Med No 83 17.7
LMTA F-DS N-DS-MS N DS-MO PIC_ Low No 52 20.9
VERT F-DS N-DS DS DS-MO PIC_ High Yes 128 14.7
  • Key to Symbols:
    B=Broad, F=Focused, N=Named, ~=Not Defined, KW=selected keywords, MS=multi synonyms,
    DS=descriptive sentence, MO=Multiple Outcomes, SC=Standard Care, EX=Exclusions, GL=Grade Level
6. Conclusion
The transformer machine learning architecture offers unique advantages for building natural language processing tools. We have found that pre-trained BERT models can be employed to generate text comparison features that outperform a variety of standard approaches for the majority of systematic review questions investigated. In addition, we observed that fine tuning one of these models often resulted in the best of all surveyed methods. However, we also found that there was no single superior approach for all research questions.
Comparison of the nature of the ten research questions in our study demonstrated that, generally, questions that had substantial detail in the natural language description of the research question performed better. The presence of explicit exclusions in the research question may also pose problems for machine learning systems. The exception to these rules was the EMR question set that had neither exclusions nor a small limited description. The EMR problem required using the larger BERT model to achieve strong performance. The EMR question statement was the second largest in the dataset and contained a complete PICO characterisation, suggesting that research questions with richer detail may require larger pre-trained models. Whether larger models, additional data or fine-tuning can transform the BERT models into consistently superior models for all use cases remains to be seen.
Acknowledgements
We thank the systematic review team ASERNIP-S (Royal Australasian College of Surgeons) for providing access to the historical set of research questions and manually reviewed article lists.
Declarations
Competing Interests
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Availability of data and materials
The dataset used for all experiments and the code used to train, test and analyse all machine learning approaches in this paper will be made available with the publication.